# Formation of extremal Reissner-Nordström black holes: insights from numerics

### v1

@@ 3.2.3 Massive scalar field
Once $\mathfrak{m}/\mathfrak{e}$ exceeds a critical value the results differ from $\mathfrak{m}=0$. With zero cosmological constant, this critical value is $0.1781$, $0.1006$ and $0.2229$ for the even, odd and polynomial Ansätze, respectively. For $\mathfrak{m}/\mathfrak{e}$ above the critical value we obtain the results shown in Figure 5 for the even Ansatz (results for the other Ansätze are qualitatively the same). For small $\mathfrak{e}M$ the results are identical to the $\mathfrak{m}=0$ results, with $\mathfrak{q}_{\rm max}$ determined by equation (21) with $\alpha_{1}=\alpha_{\rm max}$. However, once $\mathfrak{e}M$ exceeds a certain value, candidate solutions with large $\alpha_{1}$ violate the limiting condition $\partial_{U}r<0$ and so are no longer valid solutions. This leads to a kink in the plot of $\mathfrak{q}_{\rm max}$ against $\mathfrak{e}M$. This shown in more detail in the inset of Figure 5 (left). The kink occurs at lower $\mathfrak{e}M$ for larger $\mathfrak{m}/\mathfrak{e}$. Beyond the kink, $\mathfrak{q}_{\rm max}$ continues to rise, but much more slowly. For example, with the even Ansatz, when $\mathfrak{m}/\mathfrak{e}=0.2$, $\mathfrak{q}_{\rm max}$ is identical to the $\mathfrak{m}=0$ case up to the kink at $\mathfrak{e}M\approx 13.15$, but extremality is only reached much later, at $\mathfrak{e}M\approx 47.46$ (see top right panel of Figure 5), whereas the $\mathfrak{m}=0$ case reaches extremality at $\mathfrak{e}M\approx 15.96$. If $\mathfrak{m}/\mathfrak{e}$ is too large, then extremality is never attained: $\mathfrak{q}_{\rm max}$ eventually begins to decrease (extremely slowly) with larger $\mathfrak{e}M$. This is shown in the bottom right panel of Figure 5.
@@
Figure 5: Left: Maximum charge $\mathfrak{q}_{\rm max}$ as a function of $\mathfrak{e}M$ for $\mathfrak{m}/\mathfrak{e}=0$ (solid blue), $0.2$ (dashed red) and $0.5$ (dash-dotted green), with the even Ansatz. The blue curve attains extremality at $\mathfrak{e}M\approx 15.96$, denoted by a star. The latter two curves exhibit a kink when the limiting condition $\sup\partial_{U}r<0$ becomes important. The kink is more prominent for larger $\mathfrak{m}/\mathfrak{e}$; the insets show that of the red and green curves. After the kink, $\mathfrak{q}_{\rm max}$ rises very slowly with $\mathfrak{e}M$. The green curve reaches a maximum at $\mathfrak{q}<1$, and then decreases slowly. Right, top: $\mathfrak{q}_{\rm max}$ versus $\mathfrak{e}M$ with $\mathfrak{m}/\mathfrak{e}=0.2$, showing the very slow increase towards extremality, which is eventually reached at $\mathfrak{e}M\approx 47.46$, at the red star. Note the logarithmic scale on the horizontal axis. Right, bottom: The slow rise and decrease in $\mathfrak{q}_{\rm max}$ following the kink in the green curve ($\mathfrak{m}/\mathfrak{e}=0.5$). Note that all of this plot has much larger $\mathfrak{e}M$ than the kink mentioned above.
@@
This failure to attain extremality for large enough $\mathfrak{m}/\mathfrak{e}$ can be understood as follows. Notice that in (5f), the squared mass is accompanied by $|\Phi|^{2}$. In order to keep $\partial_{U}r$ negative everywhere, this term must be kept sufficiently small. The only way to keep this term small is to decrease the amplitude ($\propto\beta\equiv|\alpha_{1}|$) of the scalar field. Decreasing the amplitude requires increasing $\mathfrak{e}M$ in order to achieve, say, $\mathfrak{q}=1$. To determine the smallest value of $\mathfrak{e}M$ for which extremality is possible, we must find the largest $\beta$ such that $\sup\partial_{U}r<0$. As $\mathfrak{m}/\mathfrak{e}$ is increased, $\beta$ must decrease and $\mathfrak{e}M$ increase. Eventually, $\beta$ approaches zero, and $\mathfrak{e}M$ diverges to infinity. This implies that there is an upper bound on the value of $\mathfrak{m}/\mathfrak{e}$ for which one can reach extremality. We find that this bound is $0.2059$, $0.1114$ and $0.2773$ for the even, odd and polynomial Ansätze, respectively, with vanishing cosmological constant. These results are in agreement with [Reall:2024njy], which showed that an extremal Reissner-Nordström black hole cannot be formed in finite time in gravitational collapse if $\mathfrak{m}/\mathfrak{e}\geq 1$. [^4] This is perhaps surprising given that the proof of [Reall:2024njy] doesn’t hold for the low spacetime differentiability we are considering here. This is an example of how $C^{0}$ gluing already captures the most important features of $C^{k}$ gluing.
@@
When the cosmological constant is positive, the upper bound on $\mathfrak{m}/\mathfrak{e}$ increases slightly. For example, setting the cosmological constant to the maximum value of $\Lambda M^{2}=2/9$, with the polynomial Ansatz extremality is possible up to $\mathfrak{m}/\mathfrak{e}\approx 0.358$. This is an approximately $30\%$ rise in the upper bound compared with the asymptotically-flat case, but nevertheless remains very far from $\mathfrak{m}/\mathfrak{e}=1$. It is worth noting that reaching extremality with $\mathfrak{m}>\mathfrak{e}$ would not violate the results of [Reall:2024njy], as these apply only to asymptotically-flat spacetimes. [^5] Note also the extension to this result for supersymmetric asymptotically-AdS black holes [McSharry:2025iuz], although this doesn’t apply to the Reissner-Nordström-AdS black hole as it is not supersymmetric.
@@ 3.3.1 Restricted parameter space and $\mathfrak{Q}^{2}$
Figure 7: Left: Boundary of the restricted parameter space for the even Ansatz. Only one half is shown; the other can be obtained by a reflection through the origin. Colours indicate the value of $I$. Large values are in bluer tones, and tend to accumulate in the region where $\alpha\approx(\pm\beta,0,0)$, corresponding to a scalar field profile for which the first compactly supported pulse has much larger amplitude than the two subsequent pulses. The minimum value of $I$ on the boundary is $0.0018$ and the maximum is $0.0082$; the latter is also the global maximum. Black points denote boundary points where $\partial_{\beta}I<0$, corresponding to directions $\hat{\alpha}$ for which the maximum of II along a ray with direction $\hat{\alpha}$ is not attained at the boundary. Right: One half of the set $\mathfrak{Q}^{2}$ for the even Ansatz with $I\approx 0.0018$, the largest value such that $\mathfrak{Q}^{2}$ is topologically spherical, touching the boundary of the restricted parameter space at the North pole. The colours indicate the magnitude of $\partial_{U}\Phi(1)$ for $\mathfrak{q}=1$ and $\Lambda=\mathfrak{m}=0$, with reddish tones (located on the left side) corresponding to $\alpha$ near a candidate solution.
@@
Next we consider the surface $\mathfrak{Q}^{2}$ corresponding to parameter values that achieve the desired value of $\mathfrak{q}$ for given $\Lambda M^{2}$ and $\mathfrak{e}M$, i.e. the surface of constant $I$ defined by $\alpha$ satisfying equation (21). For small $Q/(\mathfrak{e}r_{+}^{2})$, $\mathfrak{Q}^{2}$ is a topological sphere centered at the origin [Kehle:2022uvc]. As $Q/(\mathfrak{e}r_{+}^{2})$ increases, we find that the size of the sphere increases until it touches the boundary of the restricted parameter space at the North pole (i.e. the point $\alpha=(0,0,\beta)$). This occurs for $Q/(\mathfrak{e}r_{+}^{2})\approx 0.0018$, corresponding to the surface shown on the right plot of Figure 7. At still larger $I$ a hole appears in $\mathfrak{Q}^{2}$, so its topology is no longer spherical. If $Q/(\mathfrak{e}r_{+}^{2})$ is increased even further then eventually all of $\mathfrak{Q}^{2}$ will reach the boundary and solutions of (21) will no longer exist. The qualitative behaviour of the surface $\mathfrak{Q}^{2}$ is the same for all of our Ansätze.
@@
Note that our Ansätze for $C^{0}$ gluing are not special cases of our Ansätze for $C^{1}$ gluing so we cannot compare directly our results for $C^{0}$ gluing to our results for $C^{1}$ gluing. However, we can instead use a $C^{1}$ Ansatz to perform $C^{0}$ gluing. Each point of $\mathfrak{Q}^{2}$ then corresponds to a candidate solution for $C^{0}$ gluing and the maximum value of $I$ then dictates the maximum value of $Q/(\mathfrak{e}r_{+}^{2})$ for which $C^{0}$ gluing is possible, equivalently the minimum value of $\mathfrak{e}M$ for which $C^{0}$ gluing is possible for given $\mathfrak{q}$ and $\Lambda M^{2}$. This maximum value gives $1/I_{\rm max}\approx 122$, $385$ and $4.8$ for the even, odd and polynomial Ansätze, respectively. For $\Lambda=0$ and $\mathfrak{q}=1$ this is the minimum value of $\mathfrak{e}M$ for which a candidate solution for $C^{0}$ gluing exists with this Ansatz. The limiting condition $\sup\partial_{U}r$ behaves in a qualitatively identical way to that in $C^{0}$ gluing, as in the first panel of Figure 4. In particular, whenever $\mathfrak{m}=0$, we have $\sup\partial_{U}r<0$ for all points in the restricted parameter space. Thus all points of $\mathfrak{Q}^{2}$ correspond to valid solutions of $C^{0}$ gluing.
@@ 3.3.2 Massless scalar field
All of our Ansätze exhibit qualitatively the same behaviour. For $\mathfrak{q}=1$ and $\Lambda=\mathfrak{m}=0$ we find that the minimum value of $\mathfrak{e}M$ for which a valid solution of $C^{1}$ gluing exists is $(\mathfrak{e}M)_{\rm min}=360$, $1445$, $39.3$ and $120$ for the even, odd, polynomial and modified Ansätze respectively. The optimal parameter for the modified Ansatz is now $\gamma\approx 0.46$. As would be expected, these are somewhat larger than the minimum values required when a $C^{1}$ Ansatz is used for $C^{0}$ gluing, as determined above. Notice that the polynomial Ansatz has a less constraining bound on $\mathfrak{e}M$ than the modified Ansatz, unlike in the $C^{0}$ case. This is because the latter Ansatz now is formed of three bumps, which increases the size of the derivative term $|\partial_{V}\Phi|^{2}$ in the Raychaudhuri equation (5a) but adds little charge to the system.
Figure 8: Left: Profiles $\rho$ providing solutions to the $C^{1}$ gluing problem for the even (solid black), odd (dashed blue), polynomial (dash-dotted red) and modified ($\gamma=0.46$, dotted black) Ansätze for $\mathfrak{q}=1$ and $\Lambda=\mathfrak{m}=0$. The value of $\mathfrak{e}M$ used was (approximately) the smallest that allowed $\mathfrak{q}=1$ to be reached, namely $360$, $1455$, $39.3$ and $120$, respectively. Right: Profiles for the even Ansatz with $\mathfrak{e}M=360$, for $\mathfrak{q}=0.1$, $0.6$, $0.9$, $0.99$, $0.999$ and 11 (thickest to thinnest).
@@
It is interesting to compare the minimum value of $\mathfrak{e}M$ for which $\mathfrak{Q}^{2}$ is smooth, as predicted by [Kehle:2022uvc] for large $\mathfrak{e}M$, with the (lower) minimum value of $\mathfrak{e}M$ for which a valid solution exists. This comparison gives a rough measure of how much our numerical approach goes beyond the regime of validity of the proof of [Kehle:2022uvc]. We have investigated this for the extremal case $\mathfrak{q}=1$ (again with $\Lambda=\mathfrak{m}=0$). For the even Ansatz, the former bound is $\mathfrak{e}M\gtrsim 571$, but a valid solution exists for $\mathfrak{e}M\gtrsim 360$: a decrease of nearly $37\%$. For the odd Ansatz, these bounds are $1766$ and $1455$, respectively, representing a more modest reduction of approximately $18\%$. For the polynomial Ansatz the bounds are $46.6$ and $39.3$ respectively, a decrease of $16\%$ (although the polynomial Ansatz always lies outside the regime of validity of [Kehle:2022uvc] since the latter considers only disjoint compactly-supported functions $\chi_{i}$).

### v2

@@ 3.2.3 Massive scalar field
Once $\mathfrak{m}/\mathfrak{e}$ exceeds a critical value the results differ from $\mathfrak{m}=0$. With zero cosmological constant, this critical value is $0.1876$, $0.1093$ and $0.2353$ for the even, odd and polynomial Ansätze, respectively. For $\mathfrak{m}/\mathfrak{e}$ above the critical value we obtain the results shown in Figure 5 for the even Ansatz (results for the other Ansätze are qualitatively the same). For small $\mathfrak{e}M$ the results are identical to the $\mathfrak{m}=0$ results, with $\mathfrak{q}_{\rm max}$ determined by equation (21) with $\alpha_{1}=\alpha_{\rm max}$. However, once $\mathfrak{e}M$ exceeds a certain value, candidate solutions with large $\alpha_{1}$ violate the limiting condition $\partial_{U}r<0$ and so are no longer valid solutions. This leads to a kink in the plot of $\mathfrak{q}_{\rm max}$ against $\mathfrak{e}M$. This shown in more detail in the inset of Figure 5 (left). The kink occurs at lower $\mathfrak{e}M$ for larger $\mathfrak{m}/\mathfrak{e}$. Beyond the kink, $\mathfrak{q}_{\rm max}$ continues to rise, but much more slowly. For example, with the even Ansatz, when $\mathfrak{m}/\mathfrak{e}=0.2$, $\mathfrak{q}_{\rm max}$ is identical to the $\mathfrak{m}=0$ case up to the kink at $\mathfrak{e}M\approx 14.58$, but extremality is only reached much later, at $\mathfrak{e}M\approx 25.97$ (see top right panel of Figure 5), whereas the $\mathfrak{m}=0$ case reaches extremality at $\mathfrak{e}M\approx 15.96$. If $\mathfrak{m}/\mathfrak{e}$ is too large, then extremality is never attained: $\mathfrak{q}_{\rm max}$ eventually begins to decrease (extremely slowly) with larger $\mathfrak{e}M$. This is shown in the bottom right panel of Figure 5.
@@
Figure 5: Left: Maximum charge $\mathfrak{q}_{\rm max}$ as a function of $\mathfrak{e}M$ for $\mathfrak{m}/\mathfrak{e}=0$ (solid blue), $0.2$ (dashed red) and $0.5$ (dash-dotted green), with the even Ansatz. The blue curve attains extremality at $\mathfrak{e}M\approx 15.96$, denoted by a star. The latter two curves exhibit a kink when the limiting condition $\sup\partial_{U}r<0$ becomes important. The kink is more prominent for larger $\mathfrak{m}/\mathfrak{e}$; the insets show that of the red and green curves. After the kink, $\mathfrak{q}_{\rm max}$ rises very slowly with $\mathfrak{e}M$. The green curve reaches a maximum at $\mathfrak{q}<1$, and then decreases slowly. Right, top: $\mathfrak{q}_{\rm max}$ versus $\mathfrak{e}M$ with $\mathfrak{m}/\mathfrak{e}=0.2$, showing the very slow increase towards extremality, which is eventually reached at $\mathfrak{e}M\approx 25.97$, at the red star. Note the logarithmic scale on the horizontal axis. Right, bottom: The slow rise and decrease in $\mathfrak{q}_{\rm max}$ following the kink in the green curve ($\mathfrak{m}/\mathfrak{e}=0.5$). Note that all of this plot has much larger $\mathfrak{e}M$ than the kink mentioned above.
@@
This failure to attain extremality for large enough $\mathfrak{m}/\mathfrak{e}$ can be understood as follows. Notice that in (5f), the squared mass is accompanied by $|\Phi|^{2}$. In order to keep $\partial_{U}r$ negative everywhere, this term must be kept sufficiently small. The only way to keep this term small is to decrease the amplitude ($\propto\beta\equiv|\alpha_{1}|$) of the scalar field. Decreasing the amplitude requires increasing $\mathfrak{e}M$ in order to achieve, say, $\mathfrak{q}=1$. To determine the smallest value of $\mathfrak{e}M$ for which extremality is possible, we must find the largest $\beta$ such that $\sup\partial_{U}r<0$. As $\mathfrak{m}/\mathfrak{e}$ is increased, $\beta$ must decrease and $\mathfrak{e}M$ increase. Eventually, $\beta$ approaches zero, and $\mathfrak{e}M$ diverges to infinity. This implies that there is an upper bound on the value of $\mathfrak{m}/\mathfrak{e}$ for which one can reach extremality. We find that this bound is $0.2059$, $0.1114$ and $0.2773$ for the even, odd and polynomial Ansätze, respectively, with vanishing cosmological constant. These results are in agreement with [17], which showed that an extremal Reissner-Nordström black hole cannot be formed in finite time in gravitational collapse if $\mathfrak{m}/\mathfrak{e}\geq 1$. [^4] This is perhaps surprising given that the proof of [17] doesn’t hold for the low spacetime differentiability we are considering here. This is an example of how $C^{0}$ gluing already captures the most important features of $C^{k}$ gluing.
@@
When the cosmological constant is positive, the upper bound on $\mathfrak{m}/\mathfrak{e}$ increases slightly. For example, setting the cosmological constant to the maximum value of $\Lambda M^{2}=2/9$, with the polynomial Ansatz extremality is possible up to $\mathfrak{m}/\mathfrak{e}\approx 0.3797$. This is an approximately $37\%$ rise in the upper bound compared with the asymptotically-flat case, but nevertheless remains very far from $\mathfrak{m}/\mathfrak{e}=1$. It is worth noting that reaching extremality with $\mathfrak{m}>\mathfrak{e}$ would not violate the results of [17], as these apply only to asymptotically-flat spacetimes. [^5] Note also the extension to this result for supersymmetric asymptotically-AdS black holes [16], although this doesn’t apply to the Reissner-Nordström-AdS black hole as it is not supersymmetric.
@@ 3.2.4 Modifying the Ansatz
3.2.5 Superextremal horizons
The usual definitions of extremality, sub-extremality and super-extremality only make sense for stationary spacetimes (so only for $V\geq 1$ in our situation). However, in spherical symmetry, we can formulate definitions applicable to a dynamical black hole using quasilocal mass and charge. For $\Lambda=0$ the renormalised Hawking mass $\varpi$ is defined for a symmetry sphere of area-radius $r$ by
$1-\frac{2\varpi}{r}+\frac{Q^{2}}{r^{2}}=g^{ab}(dr)_{a}(dr)_{b}$ (25)
In the Reissner-Nordström spacetime this gives $\varpi=M$ everywhere. This motivates the following definition in a dynamical spacetime, where $\varpi$ and $Q$ vary with $(U,V)$. We say that a symmetry sphere of given $(U,V)$ is subextremal if it has $\varpi>|Q|$, extremal if $\varpi=|Q|$, and superextremal if $\varpi<|Q|$. In a dynamical black hole spacetime, we can apply this definition to a symmetry sphere on the event horizon to define what we mean by the horizon being extremal, subextremal or superextremal at a given value of $V$.
For our solutions, the horizon must be sub-extremal or extremal for $V\geq 1$. However, we find that the horizon may be superextremal [^7] We thank Christoph Kehle and Ryan Unger suggesting we investigate this. for some interval (or union of intervals) of $V\in(0,1)$. It is only when the final Reissner-Nordström black hole is very far from extremality that the horizon remains subextremal for all times. For a solution that forms an extremal black hole, or a subextremal black hole of sufficiently high charge, we find that $Q>\varpi$ over a significant proportion of the full interval $V\in[0,1]$ (see Figure 7). This observation is independent of the Ansatz. [^8] Similar results hold in the higher regularity problems. At early time the horizon is always subextremal. However, often the charge increases faster, and thus eventually overtakes $\varpi$, bringing the horizon to a superextremal regime. If the Ansatz vanishes in the interior of the interval, then $\partial_{V}Q=0$ at those points, resulting in a series of “steps” in the curve of $Q$. This can lead to the horizon oscillating between subextremal and superextremal phases. An example is shown in the right panel of Figure 7. At late times, the renormalised Hawking mass increases faster than the charge, and eventually becomes equal to it (if $\mathfrak{q}=1$) or larger than it (if $\mathfrak{q}<1$). In the former case, the horizon is superextremal immediately before it becomes extremal.
Figure 7: Renormalised Hawking mass $\varpi/M$ (black) and charge $Q/M$ (dashed blue) over time. Left: Plots for a $C^{0}$ solution with the even Ansatz leading to the formation of an extremal black hole. Right: Plots for a $C^{0}$ solution with the odd Ansatz leading to the formation of a black hole with $\mathfrak{q}=0.9$. In both cases, the horizon is temporarily superextremal.
@@ 3.3.1 Restricted parameter space and $\mathfrak{Q}^{2}$
Figure 8: Left: Boundary of the restricted parameter space for the even Ansatz. Only one half is shown; the other can be obtained by a reflection through the origin. Colours indicate the value of $I$. Large values are in bluer tones, and tend to accumulate in the region where $\alpha\approx(\pm\beta,0,0)$, corresponding to a scalar field profile for which the first compactly supported pulse has much larger amplitude than the two subsequent pulses. The minimum value of $I$ on the boundary is $0.0009$ and the maximum is $0.0041$; the latter is also the global maximum. Black points denote boundary points where $\partial_{\beta}I<0$, corresponding to directions $\hat{\alpha}$ for which the maximum of II along a ray with direction $\hat{\alpha}$ is not attained at the boundary. Right: One half of the set $\mathfrak{Q}^{2}$ for the even Ansatz with $I\approx 0.0009$, the largest value such that $\mathfrak{Q}^{2}$ is topologically spherical, touching the boundary of the restricted parameter space at the North pole. The colours indicate the magnitude of $\partial_{U}\Phi(1)$ for $\mathfrak{q}=1$ and $\Lambda=\mathfrak{m}=0$, with reddish tones (located on the left side) corresponding to $\alpha$ near a candidate solution.
@@
Next we consider the surface $\mathfrak{Q}^{2}$ corresponding to parameter values that achieve the desired value of $\mathfrak{q}$ for given $\Lambda M^{2}$ and $\mathfrak{e}M$, i.e. the surface of constant $I$ defined by $\alpha$ satisfying equation (21). For small $Q/(\mathfrak{e}r_{+}^{2})$, $\mathfrak{Q}^{2}$ is a topological sphere centered at the origin [12]. As $Q/(\mathfrak{e}r_{+}^{2})$ increases, we find that the size of the sphere increases until it touches the boundary of the restricted parameter space at the North pole (i.e. the point $\alpha=(0,0,\beta)$). This occurs for $Q/(\mathfrak{e}r_{+}^{2})\approx 0.0009$, corresponding to the surface shown on the right plot of Figure 8. At still larger $I$ a hole appears in $\mathfrak{Q}^{2}$, so its topology is no longer spherical. If $Q/(\mathfrak{e}r_{+}^{2})$ is increased even further then eventually all of $\mathfrak{Q}^{2}$ will reach the boundary and solutions of (21) will no longer exist. The qualitative behaviour of the surface $\mathfrak{Q}^{2}$ is the same for all of our Ansätze.
@@
Note that our Ansätze for $C^{0}$ gluing are not special cases of our Ansätze for $C^{1}$ gluing so we cannot compare directly our results for $C^{0}$ gluing to our results for $C^{1}$ gluing. However, we can instead use a $C^{1}$ Ansatz to perform $C^{0}$ gluing. Each point of $\mathfrak{Q}^{2}$ then corresponds to a candidate solution for $C^{0}$ gluing and the maximum value of $I$ then dictates the maximum value of $Q/(\mathfrak{e}r_{+}^{2})$ for which $C^{0}$ gluing is possible, equivalently the minimum value of $\mathfrak{e}M$ for which $C^{0}$ gluing is possible for given $\mathfrak{q}$ and $\Lambda M^{2}$. This maximum value gives $1/I_{\rm max}\approx 244$, $765$ and $7.0$ for the even, odd and polynomial Ansätze, respectively. For $\Lambda=0$ and $\mathfrak{q}=1$ this is the minimum value of $\mathfrak{e}M$ for which a candidate solution for $C^{0}$ gluing exists with this Ansatz. The limiting condition $\sup\partial_{U}r$ behaves in a qualitatively identical way to that in $C^{0}$ gluing, as in the first panel of Figure 4. In particular, whenever $\mathfrak{m}=0$, we have $\sup\partial_{U}r<0$ for all points in the restricted parameter space. Thus all points of $\mathfrak{Q}^{2}$ correspond to valid solutions of $C^{0}$ gluing.
@@ 3.3.2 Massless scalar field
All of our Ansätze exhibit qualitatively the same behaviour. For $\mathfrak{q}=1$ and $\Lambda=\mathfrak{m}=0$ we find that the minimum value of $\mathfrak{e}M$ for which a valid solution of $C^{1}$ gluing exists is $(\mathfrak{e}M)_{\rm min}=720$, $2909$, $78.5$ and $239$ for the even, odd, polynomial and modified Ansätze respectively. The optimal parameter for the modified Ansatz is now $\gamma\approx 0.48$. As would be expected, these are somewhat larger than the minimum values required when a $C^{1}$ Ansatz is used for $C^{0}$ gluing, as determined above. Notice that the polynomial Ansatz has a less constraining bound on $\mathfrak{e}M$ than the modified Ansatz, unlike in the $C^{0}$ case. This is because the latter Ansatz now is formed of three bumps, which increases the size of the derivative term $|\partial_{V}\Phi|^{2}$ in the Raychaudhuri equation (5a) but adds little charge to the system.
Figure 9: Left: Profiles $\rho$ providing solutions to the $C^{1}$ gluing problem for the even (solid black), odd (dashed blue), polynomial (dash-dotted red) and modified ($\gamma=0.48$, dotted black) Ansätze for $\mathfrak{q}=1$ and $\Lambda=\mathfrak{m}=0$. The value of $\mathfrak{e}M$ used was (approximately) the smallest that allowed $\mathfrak{q}=1$ to be reached, namely $720$, $2909$, $78.5$ and $239$, respectively. Right: Profiles for the even Ansatz with $\mathfrak{e}M=720$, for $\mathfrak{q}=0.1$, $0.6$, $0.9$, $0.99$, $0.999$ and 11 (thickest to thinnest).
@@
It is interesting to compare the minimum value of $\mathfrak{e}M$ for which $\mathfrak{Q}^{2}$ is smooth, as predicted by [12] for large $\mathfrak{e}M$, with the (lower) minimum value of $\mathfrak{e}M$ for which a valid solution exists. This comparison gives a rough measure of how much our numerical approach goes beyond the regime of validity of the proof of [12]. We have investigated this for the extremal case $\mathfrak{q}=1$ (again with $\Lambda=\mathfrak{m}=0$). For the even Ansatz, the former bound is $\mathfrak{e}M\gtrsim 1141$, but a valid solution exists for $\mathfrak{e}M\gtrsim 720$: a decrease of nearly $37\%$. For the odd Ansatz, these bounds are $3531$ and $2909$, respectively, representing a more modest reduction of approximately $18\%$. For the polynomial Ansatz the bounds are $90.0$ and $78.5$ respectively, a decrease of $13\%$ (although the polynomial Ansatz always lies outside the regime of validity of [12] since the latter considers only disjoint compactly-supported functions $\chi_{i}$).